OpenStudy (anonymous):

Find the value of K so that the given point will be on the graph of the equation kx-2y=-9;(-2/3, 3)

terenzreignz (terenzreignz):

<whistles>

terenzreignz (terenzreignz):

\[\Large k\color{red}x-2\color{blue}y = -9\]

terenzreignz (terenzreignz):

Make your presence... felt :)

OpenStudy (anonymous):

:3

terenzreignz (terenzreignz):

Okay, for the point \[\Large \left(-\frac23 \ , \ 3 \right)\] To be on the line, the equation MUST be satisfied if we replace x with -2/3 and if we replace y with 3.

terenzreignz (terenzreignz):

Get it?

OpenStudy (anonymous):

oh. xD substition ??

terenzreignz (terenzreignz):

Yes... replace x with -2/3 and replace y with 3 And solve for k :) Silly ol' Derek :3

OpenStudy (anonymous):

i'm not old. -______- . :))))

terenzreignz (terenzreignz):

If you say so :3 But you do understand now, yes?

OpenStudy (anonymous):

yup.

terenzreignz (terenzreignz):

Good. That's all I need to know :)

OpenStudy (anonymous):

is the answer k=9/2??

terenzreignz (terenzreignz):

I actually don't know. I just saw this problem, knew how to solve it... but didn't actually go about solving it. Solving it now, hang on a sec...

terenzreignz (terenzreignz):

Correct.

OpenStudy (anonymous):

i have another problem which really bugs me,, i'm posting it right now.

terenzreignz (terenzreignz):

3oWhzz pFowhzzz musZtaH???

OpenStudy (anonymous):

, JEJEMON, ajajajajaja

terenzreignz (terenzreignz):

You have a problem which bugs you... how sad... <insert sad violin music>

terenzreignz (terenzreignz):

On second thought, violin music is probably too easy for you... Insert sad... trumpet music, instead, if ever there is such a piece XD

OpenStudy (anonymous):

hahaha, xD reading trumpet music is easier than violin music. xDDDDDD

OpenStudy (anonymous):

you don't use double notes in trumpet music, xDDD

terenzreignz (terenzreignz):

Of course it is... but playing a sad piece with a trumpet convincingly? mmm